Archive for the ‘Uncategorized’ Category

Broadband Problems

Posted: Wednesday, July 20th, 2016 at 09:04 by Adam Heath

We are currently aware of problems affecting the broadband platform with reports of numerous services being offline. Sources suggest this may be an issue on the BT network rather than our own, but we are continuing to investigate and gather information. Unfortunately, bt.com and btwholesale.com are also currently offline, meaning we have limited resources available to investigate at present.

We will update when further information becomes available.

  1. Adam Heath says:

    The problem has been confirmed on BT’s side and they are currently investigating. We haven’t been provided anything relating to the cause as of yet, but we will update upon receipt of this information.

    Unfortunately the BT systems remain down preventing us from looking into individual line issues.

  2. Adam Heath says:

    BT released information a short while ago advising that the problems are ‘due to power issues at one of our internet peering partners’ sites in London’. Work continues to be undertaken on BT’s side to restore service and we will update in due course.

  3. Adam Heath says:

    BT have advised that they are restoring services now. We also currently have access to some of the BT systems which were earlier unavailable. We are treating these as still at risk however until confirmation of restoration comes from BT.

  4. Adam Heath says:

    We have received confirmation from BT regarding the restoration of their systems which we access. Reports of users having no connectivity have also ceased coming into our desk, in relation to the earlier outage.

    We will be following up with BT to obtain any further formal information they release.

Telephone System

Posted: Tuesday, June 21st, 2016 at 14:57 by Ryan Berrisford

We are aware of an issue which is affecting our in-office telephone system. Inbound calls into the office are still working but some functionality of the telephone system has been lost. We are already engaged with our suppliers and further updates will follow.

  1. jonathan clarke says:

    Suppliers have advised that this issue is now a major service outage and are still working to resolve the fault as soon as possible. We will advise an update as soon as a further update is available.

  2. jonathan clarke says:

    Suppliers have advised that the issue remains at present and further monitoring and testing is taking place. Suppliers advise they will be in touch again when a further update is available.

  3. Ryan Berrisford says:

    Our suppliers have confirmed that after changes made by their Operations Team, normal service has been resumed. Furthermore our suppliers have confirmed they are carrying out investigations into the root cause of the incident.

    We apologise for any inconvenience caused to our customer base during the incident, and will post more information as it becomes available.

Incident: Virgin Media Based Circuits

Posted: Tuesday, March 31st, 2015 at 20:26 by Phil Davies

We are currently investigating an issue affecting some customers with Virgin Media provided circuits.
A fault is been raised with the supplier and we will work with them closely, until resolution. Further updates will be provided, as and when they become available.

  1. Phil Davies says:

    Our supplier has confirmed there are multiple Fibre breaks in the Gloucester area which will be affecting many of our Virgin Media provided circuits. Engineers are currently on site working on the repair, it is a major incident and we have been advised that work will continue through the night with no current estimated time of repair.

  2. Phil Davies says:

    Our Supplier has informed us that work continued through the night with most circuits having a restored service at 22:30 due to re-routing within their network however the physical work to is due to complete around 10.30 today so further downtime may occur.

Incident: Telehouse Pwan Switch Blade Failure

Posted: Monday, February 2nd, 2015 at 04:10 by Ming-Yu Hsieh

We are currently investigating what appears to be a blade failure within our telehouse pwan switch router.

Further updates will be provided as and when they become available.

  1. Ming-Yu Hsieh says:

    Entanet engineers have been dispatched to telehouse with replacement hardware. We will update again once our engineers have arrived..

  2. Wayne Hawtin says:

    The latest update is that engineers are still en-route with an ETA of 09:30 due to traffic. Further updates will be provided when they become available.

  3. Wayne Hawtin says:

    We have had confirmation that the engineers have now arrived onsite. Further updates will follow as they become available.

  4. Steve Lalonde says:

    The faulty line card has been replaced, all services are restored.

    Please contact support if your service is still affected.

  5. Wayne Hawtin says:

    An incident report is now available and can be found here.

Planned work: Interxion DSL power feeds

Posted: Friday, August 22nd, 2014 at 14:49 by David Derrick

Thu 4 Sep 22:00 to Fri 5 Sep 02:00 BST

During the above maintenance window we will be moving all single fed DSL equipment in Interxion onto a new dual fed power supply. This will involve power cycling a number of LNS and LTS routers which will disconnect any users connected at the time. Sessions will only reconnect to a device which is already on dual power so users should not experience multiple drops.

Users experiencing connectivity issues after the maintenance window has closed are advised to reboot their routers before contacting the support desk.

  1. Steve Lalonde says:

    This work is now complete.

Planned Maintenance: Server Relocation

Posted: Wednesday, July 23rd, 2014 at 12:28 by Ming-Yu Hsieh

00:01 – 06:00 Tuesday 29th July 2014

During the above window we will be continuing to relocate internal servers within our Telford datacentres.

We expect downtime of approximately 30 mins during this window.

We apologise for any inconvenience this may cause.

  1. Ming-Yu Hsieh says:

    This work is now complete.

21C Connectivity Issue

Posted: Monday, June 16th, 2014 at 11:40 by Neil Watson

We are aware of an issue affecting 21CN connectivity in the Norfolk, Suffolk and Essex areas. Affected services appear to be both 21C Broadband and Ethernet services, including EFM type circuits.

For DSL services, the typical symptoms appear to be prolonged periods without sync with brief periods of connectivity in between.

Exchanges where we have received outage confirmation from BT are:

Norwich
St Osyth
Cromer
Holbrook
Dedham
Harwich
Gorleston
Felixstowe
Little Clacton
Kesgrave
Earls Colne
Rattlesden
Sawston
Great Wenham
Bury St Edmunds
Great Horkesley
Kentford

Ethernet based services are intermittently losing connection.

BT are aware of the issue and we’re working with them to ensure the soonest possible resolution.

  1. Ryan Berrisford says:

    Full service has been restored to the affected customers and we are seeing stability since approximately 11:25.

    BT have confirmed that the incident was caused by a configuration issue which has now been rectified.

Incident: telford-dc1.core.enta.net

Posted: Sunday, May 11th, 2014 at 00:24 by Ming-Yu Hsieh

We are currently aware of an issue affecting connectivity to our telford-dc1.core router. Engineers are working on the issue and further updates will follow.

Anything directly connected to the telford-dc1.core will be affected.

  1. Ming-Yu Hsieh says:

    We are currently replacing faulty line card.

  2. Ming-Yu Hsieh says:

    All service has been restored.

Customer DDOS attack

Posted: Monday, February 17th, 2014 at 13:04 by Neil Watson

We are aware that an individual customer has been the victim of a large DDOS attack which has had some impact on our network. During the attack performance for some customers may have been degraded. We were able to identify the affected customer and took corrective action to remove the traffic from our network. Levels have now returned to normal and we will continue to monitor for further attacks.

Incident: xDSL connectivity

Posted: Friday, January 17th, 2014 at 16:40 by Richard Partridge

We are aware of an issue currently affecting our xDSL customer base, manifesting itself as poor performance.

Our Systems team are currently investigating. Further information will be provided as and when it becomes available.

  1. Richard Partridge says:

    Throughput levels across the broadband network have now returned to normal. We are currently monitoring the situation closely whilst our investigation continues.

  2. Richard Partridge says:

    We’ve identified a residual issue affecting EWCS customers. A work around is currently being implemented.

    Further work may be required to resolve the issue. More information on this will be provided once we finalise a plan of action.

  3. Richard Partridge says:

    The work around has now been deployed and we’re seeing improvement in bandwidth utilisation for EWCS based customers.

    Further updates will follow.

  4. Steve Lalonde says:

    Subsequent engineering work has resolved the issues and the network is now performing correctly.

    An investigation is being conducted into the initial cause of the fault and an incident report will be made available next week.

  5. Richard Partridge says:

    An incident report has been made available for this issue. It can be found here.